# ThinkGeo [Uncaptioned image] : Evaluating Tool-Augmented Agents for Remote Sensing Tasks

### v1

@@ 4 Tool Suite and Evaluation
Task Format: Each task is posed as a step and tool-implicit query, requiring the agent to reason and respond in a ReAct-style format [35]. Agents autonomously generate thought steps, select tools from a predefined set, format arguments, and produce final answers, evaluating spatial reasoning, planning, and multi-step execution grounded in remote sensing imagery.
Tool Categories: ThinkGeo extends the AgentLego framework [3] by integrating two additional tools: ChangeDetection [14] for multi-temporal remote sensing analysis and SegmentObjectPixels [17, 18] for segmentation and get the segmented object pixels count. The toolset is organized into three functional categories: Perception (e.g., TextToBbox, CountGivenObject, ChangeDetection), Logic, and Operation, supporting object localization, spatial reasoning, and interactive annotation. Logic tools (e.g., Calculator, Solver) support numerical reasoning, distance calculations, and spatial comparisons. Operation tools (e.g., DrawBox, AddText, GoogleSearch) facilitate visual annotation and output formatting. This categorization supports both fine-grained evaluation (e.g., tool-category performance) and structured analysis of agent planning behavior across spatial, logical, and domain-specific subtasks.
Evaluation Methodology: We adopt the evaluation framework of GTA [28] for step-by-step metrics, including instruction-following (InstAcc), tool selection (ToolAcc), argument correctness (ArgAcc), and summary generation (SummAcc), to enable fine-grained assessment of agent behavior. However, GTA computes final answer accuracy (AnsAcc) using deterministic string matching, which can misclassify predictions as correct & incorrect due to minor phrasing variations. To mitigate this, we introduce LLM-as-a-judge for evaluating final answers: for each query, we construct curated evaluation questions and use 4o-mini to assess whether the model’s prediction satisfies the correct answer. This offers a more reliable measure of task success, especially for multi-fact answers. All other metrics and tool-based evaluations follow the GTA protocol.
@@
Table 3: Evaluation results across models on the ThinkGeo benchmark. The table reports step-by-step execution metrics (left) and end-to-end evaluation results (right), including tool-type accuracy (P: Perception, O: Operation, L: Logic), Ans. (final answer), and answer accuracy under image grounding (Ans_I). Overall, GPT4 family performs the best.
Model Step-by-Step Metrics End-to-End Metrics
Inst. Tool. Arg. Summ. P. O. L. Ans. Ans_I
GPT-4o 73.31 63.75 33.31 59.15 87.05 76.68 67.88 11.51 20.02
GPT-4-1106 82.44 73.21 37.74 64.22 79.91 69.15 56.29 9.46 16.91
Qwen1.5-7b-chat 23.44 8.67 3.26 45.17 4.94 37.25 20.98 6.91 6.74
Yi-1.5-6B-Chat 22.40 26.47 0.26 32.16 5.56 2.22 4.64 9.97 5.83
5 Experiments & Discussion
To assess the reasoning and tool-use capabilities of language models under real-world remote sensing scenarios, we conduct comprehensive evaluations on the ThinkGeo benchmark. Our benchmark poses multimodal and tool-implicit challenges that require agentic models to invoke tools across perception, operation, and logic categories. Unlike prior evaluations that rely on synthetic queries or shallow tool interactions, our benchmark emphasizes realistic queries grounded in satellite or aerial imagery and demands multi-step reasoning with spatial and numerical precision.
Quantitative Analysis: We evaluate a wide range of models, including GPT-4o [13], GPT-4-1106 [2], and several open-source variants (e.g., Qwen [12, 4], InternLM [6], LLaMA3 [8], Phi [1], and Mistral [15]), in both step-by-step and end-to-end settings. The step-by-step mode evaluates intermediate stages such as instruction following (Inst.), tool selection (Tool.), argument formatting (Arg.), and summary generation (Summ.). The end-to-end mode measures performance on tool categories (P: Perception, O: Operation, L: Logic), final answer correctness (Ans.), and answer correctness under visual grounding (Ans_I). As shown in Table 3, GPT-4o and GPT-4-1106 achieve the relatively better overall accuracy, showcasing superior planning and execution across multi-step tool chains. However, most open-source models struggle with tool call formatting and argument prediction, resulting in significantly lower answer accuracy. Among all step-by-step metrics, the tool has more correlation with final answer accuracy, underscoring its importance in agentic performance (Fig. 5).
Tool Call & Error: The distribution of tool calls and error rates highlights key differences in agentic reliability across models. Proprietary models like GPT-4o and GPT-4-1106 show frequent tool use with relatively low error rates (47.3% and 30.0%), indicating strong tool-handling capabilities (Fig. 5). In contrast, open-source models such as Qwen3-8B, InternLM3-8B, and LLaMA3-8B invoke tools aggressively but incur high error rates, reflecting poor execution control. Meanwhile, smaller models like Qwen1.5-7B and Phi-3 often fail despite limited tool use, underscoring challenges in formatting and context alignment. These trends suggest that effective agent behavior hinges not just on tool access but on precise invocation and robust reasoning.
Easy vs Hard Queries: We analyze the performance of LLM agents on queries of varying difficulty levels, as defined in Section 3.3, which includes 281 easy and 155 hard queries. Figure 6 presents a bar chart showing the number of correct responses per model, separated by difficulty level. The x-axis lists the evaluated models, while the y-axis indicates the count of correctly answered queries. Blue bars represent easy queries, and red bars denote hard ones. This analysis highlights a consistent performance gap across difficulty levels, emphasizing the increased challenges LLMs face when dealing with complex, multi-step reasoning tasks.
6 Conclusion
We propose ThinkGeo, the first benchmark tailored specifically to evaluate tool-augmented LLM agents on real-world RS tasks. Since ThinkGeo grounds evaluation in high-resolution EO imagery, structured tool-use pipelines, and fine-grained reasoning annotations, it reveals critical gaps in current agent capabilities. In particular, our analysis shows room for improvement in spatial planning, temporal consistency, and domain-specific tool integration. Our extensive study across 436 tasks and multiple SoTA LLMs demonstrates that, while tools like segmentation and change-detection improve raw perception, true geospatial reasoning remains an open challenge. ThinkGeo aims to attract further efforts towards the development of next-generation multimodal agents that can seamlessly blend perception, planning, and execution in complex, spatially grounded RS and EO environments.
Limitations: While ThinkGeo covers a diverse set of RS scenarios, it focuses mainly on optical imagery. Other modalities (e.g., SAR, LiDAR) and more specialized geospatial operations are not yet represented in the benchmark. Although our task set provides broad coverage, further scaling and community-driven extensions will be necessary to fully capture the long tail of geospatial use cases.
Supplementary Material
This supplementary material provides extended analysis and additional results to support the main paper. It includes evaluations on the Claude model (S1), and detailed error breakdowns (S2). We also present representative samples across use case categories (S3), tool usage distribution (S4), and category-wise sample counts (S5). Additionally, we report the human effort (S6) involved in curating, generating, and verifying the proposed ThinkGeo benchmark.
Appendix S1 Claude Model Evaluation
As shown in the main paper, GPT-4 models overall perform well across both planning and execution stages. In comparison, Claude’s results suggest it can follow through effectively once the right tools are selected, but often fails to choose or use them correctly on its own. Specifically, Claude-3.7-Sonnet shows better performance in usage of operational tools and summarization of final result, but struggles with tool selection (26.21%) and argument prediction (0.33%), which limits its reliability in multi-step tasks (Tab. A1). With the inclusion of Claude-3.7-Sonnet, the updated tool call and error is shown in Fig. A1.
Table A1: Evaluation results across models on the ThinkGeo benchmark. The table reports step-by-step execution metrics (left) and end-to-end evaluation results (right), including tool-type accuracy (P: Perception, O: Operation, L: Logic), Ans. (final answer), and answer accuracy under image grounding (Ans_I). Overall, GPT4 family performs the best.
@@ 6 Conclusion
Figure A1: The plot illustrates the total number of tool calls made by each model and the corresponding number of tool call errors. The large discrepancy in open source models indicates a high rate of tool misuse. In contrast, models like GPT-4o demonstrate better tool invocation reliability.
Appendix S2 Error Analysis
@@
The error analysis summarized in Tab. A2 provides insights into the types of prediction failures exhibited by different models on the ThinkGeo benchmark, categorized into planning and format-related errors. Planning errors, specifically NoAction, is prominent for models like GPT-4-1106 (84.87%), GPT-4o (95.82%), and Qwen1.5-7b (94.90%), indicating extra reasoning or summaries without delivering a final actionable response, i.e., either a tool call or an explicit answer. In contrast, models such as LLama3-1-8b and Qwen2.5-7b show high rates of "Invalid JSON" errors (82.89% and 84.44%, respectively), revealing difficulty in producing syntactically correct tool input formats even when the correct tool is selected. Among format-related errors, "Final Answer (SingleStep)" exhibits high rates: Phi-3-mini-4k (73.74%), Qwen1.5-7b (80.05%), and Qwen2.5-7b (74.94%), where models bypass intermediate reasoning and prematurely generate a final answer. In particular, GPT-4o and GPT-4-1106 show remarkably low error rates in the format category, demonstrating strong capabilities in structured reasoning. Overall, the results highlight that models face different challenges, some struggle with planning and step-by-step reasoning, while others have issues formatting tool inputs or following the required response structure.
@@
Table A2: Breakdown of errors made by models on the ThinkGeo benchmark. Errors are grouped into planning and format-related categories (percentages are reported).
@@
Appendix S3 More Samples by Use Case
Fig. A2 shows examples in which ThinkGeo queries prompt-driven agents to compose multiple tools from the available set. Each ReAct-style execution chain demonstrates spatial reasoning and multi-step decision making grounded in satellite imagery, reflecting the benchmark’s focus on real, tool-oriented geospatial problem solving.
Figure A2: Representative examples from the ThinkGeo benchmark. Each row shows a user query (left), the corresponding ReAct-style execution chain involving tool calls (center), and the final answer (right).
Appendix S4 Tool Usage Distribution
The distribution of tool usage across the ThinkGeo benchmark is shown in Fig. A3. The most frequently invoked tools are Calculator, TextToBbox, and RegionAttributeDescription, reflecting the benchmark’s emphasis on spatial computation, object localization, and attribute reasoning. Mid-frequency tools such as CountGivenObject and ChangeDetection support core analysis tasks, while tools such as AddText, ObjectDetection, and OCR are rarely used, indicating their narrower application scope. This spread highlights the diversity in tool reliance and the complexity of multi-step reasoning across geospatial scenarios.
Figure A3: Tool-wise usage distribution in the ThinkGeo benchmark.
Appendix S5 Category-wise Sample Statistics
Tab. A3 summarizes the ThinkGeo benchmark by task category and difficulty level. Among the seven core domains, Disaster Assessment & Change Analysis contributes the largest share (148 queries), reflecting the complexity of temporal reasoning and damage interpretation. Urban Planning and Transportation Analysis are followed by diverse spatial reasoning tasks.
Table A3: Category-wise count of easy and hard tasks, including totals.
@@
Appendix S6 Human Annotation Effort
[…]

### v2

@@ 4 Tool Suite and Evaluation
Task Format: Each task is posed as a step and tool-implicit query, requiring the agent to reason and respond in a ReAct-style format Yao et al. (2023). Agents autonomously generate thought steps, select tools from a predefined set, format arguments, and produce final answers, evaluating spatial reasoning, planning, and multi-step execution grounded in remote sensing imagery.
@@
Table 3: Evaluation results across models on the ThinkGeo benchmark. The table reports step-by-step (left) and end-to-end evaluation results (right), including tool-type accuracy (P: Perception, O: Operation, L: Logic), Ans. (final answer), and answer accuracy under image grounding (Ans_I). Overall, GPT4 family performs the best.
@@
Tool Categories: ThinkGeo extends the AgentLego framework AgentLego Contributors (2023) with two additional tools: ChangeDetection Irvin et al. (2024) for multi-temporal remote sensing analysis and SegmentObjectPixels Kirillov et al. (2023); Li et al. (2022) for segmentation and pixels counting. The toolset is organized into three functional categories: Perception (e.g., TextToBbox, ChangeDetection), Logic, and Operation, supporting object localization, spatial reasoning, and interactive annotation. Logic tools (e.g., Calculator, Solver) support numerical reasoning, distance calculations, and spatial comparisons. Operation tools (e.g., DrawBox, GoogleSearch) facilitate visual annotation and output formatting. This categorization supports fine-grained evaluation (e.g., tool-category performance) and structured analysis of planning behavior across spatial, logical, and domain-specific subtasks.
Evaluation Methodology: We adopt the evaluation framework of GTA Wang et al. (2024) for step-by-step metrics, including instruction-following (InstAcc), tool selection (ToolAcc), argument correctness (ArgAcc), and summary generation (SummAcc), to assess agent behavior. While GTA computes final answer accuracy (AnsAcc) using deterministic string matching, this can misclassify predictions due to variations in phrases. To mitigate this, we introduce LLM-as-a-judge: curated evaluation questions per query and use 4o-mini to verify the correctness of the model’s prediction. This offers a more reliable measure of task success, especially for multi-fact answers.
5 Experiments & Discussion
To assess the reasoning and tool-use capabilities of language models under real-world remote sensing scenarios, we conduct comprehensive evaluations on the ThinkGeo benchmark.
Our benchmark poses multimodal and tool-implicit challenges that require agentic models to invoke tools across perception, operation, and logic categories. Unlike prior evaluations that rely on synthetic queries or shallow tool interactions, our benchmark emphasizes realistic queries grounded in satellite or aerial imagery and demands multi-step reasoning with spatial and numerical precision.
Quantitative Analysis: We evaluate a wide range of models, including GPT-4o Hurst et al. (2024), GPT-4-1106 Achiam et al. (2023), and several open-source variants (e.g., Qwen Hui et al. (2024); Bai et al. (2023), InternLM Cai et al. (2024), LLaMA3 Grattafiori et al. (2024), Phi Abdin et al. (2024), and Mistral Jiang et al. (2023)), in both step-by-step and end-to-end settings. The step-by-step mode evaluates intermediate stages such as instruction following (Inst.), tool selection (Tool.), argument formatting (Arg.), and summary generation (Summ.). The end-to-end mode measures performance on tool categories (P: Perception, O: Operation, L: Logic), final answer correctness (Ans.), and answer correctness under visual grounding (Ans_I). All evaluations and analyses presented in the paper are conducted on optical RGB imagery, ensuring consistency and comparability across models. To demonstrate the extension of our framework to additional modalities, we present SAR-based analysis in Appendix (section  A.1). As reported in Table 3, GPT-4o and GPT-4-1106 achieve the strongest overall accuracy, reflecting superior planning and execution across multi-step tool chains. Most open-source models struggle with tool call formatting and argument prediction, resulting in significantly lower accuracy of the answers. Among all step-by-step metrics, tool selection has the highest correlation with final answer accuracy, underscoring its importance in agentic performance (Fig. 6).
Tool Call & Error: Tool calls and error rates highlight key gaps in agentic reliability. Proprietary models (GPT-4o, GPT-4-1106) show frequent tool use with relatively low error rates (47.3% and 30.0%), indicating strong tool-handling capabilities (Fig. 6). In contrast, open-source models (Qwen3-8B, InternLM3-8B, LLaMA3-8B) invoke tools aggressively but incur high error rates, reflecting poor execution control. Meanwhile, smaller models ( Qwen1.5-7B, Phi-3) often fail despite limited tool use, underscoring formatting and context alignment issues. These trends suggest that effective agent behavior hinges not just on tool access but on precise invocation and robust reasoning.
Easy vs Hard Queries: We analyze the performance of LLM agents on queries of varying difficulty levels, as defined in Section 3.3 Figure 4 presents a bar chart of correct responses per model, separated by difficulty level. The x-axis lists the evaluated models, while the y-axis indicates the count of correctly answered queries. Blue bars represent easy queries, and red bars denote hard ones. This analysis highlights a consistent performance gap across difficulty levels, emphasizing the increased challenges LLMs face when dealing with complex, multi-step reasoning tasks.
6 Conclusion
We propose ThinkGeo, the first benchmark tailored specifically to evaluate tool-augmented LLM agents on real-world RS tasks. Since ThinkGeo grounds evaluation in high-resolution EO imagery, structured tool-use pipelines, and fine-grained reasoning annotations, it reveals critical gaps in current agent capabilities. In particular, our analysis shows room for improvement in spatial planning, temporal consistency, and domain-specific tool integration. Our extensive study across 486 tasks and multiple SoTA LLMs demonstrates that, while tools like segmentation and change-detection improve raw perception, true geospatial reasoning remains an open challenge. ThinkGeo aims to attract further efforts towards the development of next-generation multimodal agents that can seamlessly blend perception, planning, and execution in complex, spatially grounded RS and EO environments.
Appendix A Appendix
This supplementary material provides extended analysis and additional results to support the main paper. It includes evaluations on the SAR data(A.1), and detailed error breakdowns (A.2). We also present representative samples across use case categories (A.3), tool usage distribution (A.4), and category-wise sample counts (A.5). Moreover, we report the human effort (A.6) involved in curating, generating, and verifying the proposed ThinkGeo benchmark. Additionally, we analyze the runtime performance of the evaluated models (A.7)
A.1 SAR Data Evaluation
A set of 50 queries over 244 tool-reasoning steps utilize SAR imagery. This data is developed through the same rigorous, three-phase manual curation pipeline described above, ensuring identical depth and consistency in annotation; evaluation results are reported in Table 4. Proprietary models like GPT-4 variants lead in instruction and tool usage accuracy on SAR imagery, showing strong generalization in structured reasoning. However, final answer accuracy remains low ($10\%$) across all models, revealing that even the best proprietary models struggle with visual grounding in non-optical scenarios. This highlights a key gap: tool competence alone is not sufficient for reliable SAR understanding.
Table 4: Evaluation results across models on the ThinkGeo SAR benchmark. The table reports step-by-step execution metrics (left) and end-to-end evaluation results (right), including tool-type accuracy (P: Perception, O: Operation, L: Logic), Ans. (final answer), and answer accuracy under image grounding (Ans_I).
Model Step-by-Step Metrics End-to-End Metrics
Inst. Tool. Arg. Summ. P. O. L. Ans. Ans_I
GPT-4o 75.00 69.67 49.18 50.00 87.43 81.82 48.15 5.56 22.29
GPT-4-1106 75.71 71.31 48.77 58.33 56.93 70.45 17.45 5.56 17.10
Claude-3.7-Sonnet 14.64 21.72 0.41 77.78 85.93 55.56 70.06 8.33 24.00
Qwen1.5-7b-chat 15.36 4.10 2.05 36.11 0.00 47.06 1.44 2.78 10.00
Qwen2.5-7b-Instruct 56.43 44.26 29.10 61.11 16.49 26.67 27.97 5.56 8.00
InternLM3-8b-Instruct 51.79 47.13 29.92 36.11 30.87 30.51 29.81 0.00 7.74
LLaMA3-1-8b-Instruct 40.71 34.43 20.08 55.56 56.38 25.35 46.06 2.78 8.00
Phi-3-mini-4k-Instruct 33.21 29.51 16.80 33.33 21.05 25.00 24.52 5.56 8.00
A.2 Error Analysis
@@
The error analysis summarized in Tab. 5 provides insights into the types of prediction failures exhibited by different models on the ThinkGeo benchmark, categorized into planning and format-related errors. Planning errors, specifically NoAction, is prominent for models like GPT-4-1106 (84.87%), GPT-4o (95.82%), and Qwen1.5-7b (94.90%), indicating extra reasoning or summaries without delivering a final actionable response, i.e., either a tool call or an explicit answer. In contrast, models such as LLama3-1-8b and Qwen2.5-7b show high rates of ”Invalid JSON” errors (82.89% and 84.44%, respectively), revealing difficulty in producing syntactically correct tool input formats even when the correct tool is selected. Among format-related errors, ”Final Answer (SingleStep)” exhibits high rates: Phi-3-mini-4k (73.74%), Qwen1.5-7b (80.05%), and Qwen2.5-7b (74.94%), where models bypass intermediate reasoning and prematurely generate a final answer. In particular, GPT-4o and GPT-4-1106 show remarkably low error rates in the format category, demonstrating strong capabilities in structured reasoning. Overall, the results highlight that models face different challenges, some struggle with planning and step-by-step reasoning, while others have issues formatting tool inputs or following the required response structure.
@@
Table 5: Breakdown of errors made by models on the ThinkGeo benchmark. Errors are grouped into planning and format-related categories (percentages are reported).
@@
A.3 More Samples by Use Case
Fig. 8 shows examples in which ThinkGeo queries prompt-driven agents to compose multiple tools from the available set. Each ReAct-style execution chain demonstrates spatial reasoning and multi-step decision making grounded in satellite imagery, reflecting the benchmark’s focus on real, tool-oriented geospatial problem solving. Additionally, we illustrate complete reasoning trajectories and grounding in queries in Fig. 10 and  11
Figure 8: Representative examples from the ThinkGeo benchmark. Each row shows a user query (left), the corresponding ReAct-style execution chain involving tool calls (center), and the final answer (right).
A.4 Tool Usage Distribution
The distribution of tool usage across the ThinkGeo benchmark is shown in Fig. 9. The most frequently invoked tools are Calculator, TextToBbox, and RegionAttributeDescription, reflecting the benchmark’s emphasis on spatial computation, object localization, and attribute reasoning. Mid-frequency tools such as CountGivenObject and ChangeDetection support core analysis tasks, while tools such as AddText, ObjectDetection, and OCR are rarely used, indicating their narrower application scope. This spread highlights the diversity in tool reliance and the complexity of multi-step reasoning across geospatial scenarios.
Figure 9: Tool-wise usage distribution in the ThinkGeo benchmark.
A.5 Category-wise Sample Statistics
Tab. 6 summarizes the ThinkGeo benchmark by task category and difficulty level. Among the seven core domains, Disaster Assessment & Change Analysis contributes the largest share (148 queries), reflecting the complexity of temporal reasoning and damage interpretation. Urban Planning and Transportation Analysis are followed by diverse spatial reasoning tasks.
Table 6: Category-wise count of easy and hard tasks, including totals.
@@ A.5 Category-wise Sample Statistics
A.6 Human Annotation Effort
The construction of the ThinkGeo benchmark involved a coordinated annotation effort totaling $\approx 500$ human hours, distributed across a 4-member team. This time was spent across three core stages:
@@ A.6 Human Annotation Effort
Figure 10: Example of a tool-augmented reasoning in ThinkGeo benchmark query. The illustration shows a multi-step task where the agent is required to combine perception, logic, and spatial reasoning tools in sequence. The figure highlights how queries are grounded in real remote sensing imagery and require tool-augmented reasoning chains to arrive at correct geospatial conclusions.
A.7 Runtime Analysis
To complement the evaluation results, we analyze the runtime performance of the evaluated models. Table 7 reports both the step-by-step latency (measuring average runtime per query) and the end-to-end pipeline latency (including LLM inference, tool calls, and result aggregation). These results highlight the computational cost of agentic reasoning for geospatial tasks.
Table 7: Average runtime per query across evaluated models. Step-by-step latency corresponds to the time required to execute all reasoning steps; end-to-end latency accounts for the entire pipeline execution.
Model Step-by-step Avg. (s/query) End-to-end Avg. (s/query)
GPT-4o 12.10 12.01
GPT-4-1106 22.43 16.65
Qwen1.5-7b-chat 22.54 6.79
Qwen2.5-7B-Instruct 16.79 7.17
InternLM3-8b-Instruct 22.68 27.01
LLaMA3-1-8b-Instruct 13.16 14.78
While proprietary models such as GPT-4o and GPT-4-1106 offer relatively stable runtime profiles, open-source models often display higher variability across step-by-step and end-to-end latencies. Notably, InternLM3-8b exhibits significant end-to-end overhead. Step-by-step evaluation is longer for some models, like Qwen1.5, despite appearing more incremental in nature. These models process each action prediction by re-encoding the entire history of previous steps as context. This cumulative context grows with each step, significantly increasing the load and inference time. By contrast, in the end-to-end setting, Qwen1.5 often generates answers with fewer tool calls or shorter reasoning traces, resulting in less overall computation.
These findings align with prior observations that current agentic pipelines continue to incur substantial latency Mialon et al. (2023). Reducing end-to-end response time remains an important open research direction for enabling real-time geospatial applications.
Figure 11: Example of change detection in ThinkGeo Post-disaster damage assessment and resource planning. The figure illustrates a task that requires the identification and comparison of temporal differences between pre-disaster and post-disaster imagery. The ReAct-style execution chain illustrates how agents must invoke perception and computation tools to quantify changes, such as structural damage.
LLM Usage Statement: We used large language models exclusively for polishing the writing and providing assistance during dataset curation. They were not involved in research ideation, model design, and analysis.